OpenStudy (anonymous):

Which two numbers below are equal? [Select two of the numbers below.] one and one over five one and three over fifteen three and one over fifteen thirteen over fifteen

OpenStudy (anonymous):

Hint 18 = 15 + 3

OpenStudy (anonymous):

@pgpilot326 thats a. and I need one more...

OpenStudy (anonymous):

\[\frac{ 18 }{15 }=\frac{ 15+3 }{ 15 }=\frac{ 15 }{ 15 }+\frac{ 3 }{ 15 }=?\]

OpenStudy (anonymous):

@pgpilot326 a and d

OpenStudy (anonymous):

Well you have: \[\frac{ 18 }{ 15 } = \frac{ 15 }{ 15 }+\frac{ 3 }{ 15 }\]\[1+\frac{ 3 }{ 15 }=\frac{ 15 }{ 15 }+\frac{ 3 }{ 15 }\]\[3+\frac{ 1 }{ 15 }=\frac{ 3*15 }{ 15 }+\frac{ 1 }{ 15 }\]\[\frac{ 13 }{ 15 }=\frac{ 13 }{ 15 }\] which do you think are the same?

OpenStudy (anonymous):

I think its... \[\frac{ 18 }{ 15 } and \frac{ 3 }{ 15 }\]

OpenStudy (anonymous):

@sarahusher

OpenStudy (anonymous):

Okay 3/15 is not one of your options 18/15 is correct, but look at the sums I did above, which of the sums give you 18/15?

OpenStudy (anonymous):

@sarahusher there was a typo with the copy and paste and a wasnt 18/15 it was 1 1/5 so I fixed it you can look at the problem again

OpenStudy (anonymous):

@sarahusher hello?

OpenStudy (anonymous):

Okay :) no problem, so it's: A) one and one over five \[1+\frac{ 1 }{ 5 }=\frac{ 5 }{ 5 }+\frac{ 1 }{ 5 }=\frac{ 6 }{ 5 }\] B) one and three over fifteen \[1+\frac{ 3 }{ 15 }=\frac{ 15 }{ 15 }+\frac{ 3 }{ 15 }=\frac{ 18 }{ 15 }=\frac{ 6 }{ 5 }\] C) three and one over fifteen \[3+\frac{ 1 }{ 15 }=\frac{ 3*15 }{ 15 }+\frac{ 1 }{ 15} =\frac{ 46 }{ 15 } \] D) thirteen over fifteen \[\frac{ 13 }{ 15 }\] Looking at those, which two are the same?

OpenStudy (anonymous):

a and d still right lol I thought that

OpenStudy (anonymous):

or a and b

OpenStudy (anonymous):

If you write the answers A and B A and D next to each other, which are the same? :)